# Towards a Mechanistic Explanation of Diffusion Model Generalization

### v2

@@ 5.1 PSPC-Square
One reasonable patch sets is $\mathcal{C}s$, the set of overlapping square patches of spatial size $s\times s$ defined in Section 4. However, as shown by Figure 4, optimal patch size varies substantially across the reverse diffusion process. To adapt our method to this observation, we define a patch size schedule $s(t):\mathbb{R}^{+}\rightarrow\{1,\ldots,h\}$. The square Patch Set Posterior Composite (PSPC-Square) is defined by Equation 9 with $\mathcal{C}=\mathcal{C}_{s(t)}$. In practice, we set s(t) according to Figure 4, choosing patch size at each $t$ with the lowest error vs. the network denoiser. Patch size schedules can be found in Appendix B.
@@ 5.2 PSPC-Flex
While square patches mimic the square receptive fields of convolutional U-Nets, Figure 3 suggests that denoisers gradients are not perfectly square. To enable more flexible, non-square patches of varying shapes and sizes, we introduce PSPC-Flex, which utilizes an adaptive patch set based on average gradient sensitivity maps.
@@
Figure 8: PF-ODE sampling trajectories starting from shared initial $\mathbf{z}$ using a DDPM++ score estimator (Left) or our PSPC-Flex estimator (Right). At each sigma, we display the noisy image $\mathbf{z}$ along with the output of the DDPM++ denoiser and PSPC-Square on that $\mathbf{z}$. In all cases, minor differences between DDPM++ and PSPC-Square outputs only occur over intermediate $t$. Sampling using PSPC-Flex yields samples which are similar in content to the network samples. However, compounding errors lead to suboptimal final samples.
@@ 6.1 Network Denoiser Comparison
To evaluate the similarity of PSPC to network denoisers, we compare various empirical denoisers against the EDM DDPM++ denoiser outputs for each dataset. In addition to our PSPC-Square an PSPC-Flex methods, we evaluate the empirical denoiser of Equation 7, and the Gaussian denoiser decribed by Wang & Vastola (2024); Li et al. (2024).
Figure 6 plots the MSE of each of the denoisers against the neural network denoiser. We find that for both forward and reverse processes, both PSPC variants have substantially lower MSE against the network denoiser outputs as compared to the empirical denoiser. Compared to the Gaussian denoiser, our methods generally have better MSE for $t\in[0.3,3]$. Comparing our methods, the flexibility of the PSPC-Flex patch set produces better estimates than those of PSPC-Square. Over the reverse process evaluation sets on every dataset, PSPC-Flex generally estimates the network denoiser better than PSPC-Square and all other methods.
Qualitatively, we compare the output of PSPC-Flex to several CIFAR-10 network denoisers and the empirical denoiser in Figure 1. For every $t$, PSPC-Flex outputs are visually similar to the outputs of the various network denoisers, suggesting that local denoising operations comprise a significant portion of the generalization mechanism of diffusion models. Additional denoiser outputs can be found in Appendix D
@@ 6.2 PSPC Sampling
Inspired by the remarkable similarity between network denoisers and PSPC-Flex evidenced by Figures 1 and 6, we investigate the efficacy of PSPC-Flex as fully training-free, completely non-neural diffusion model.
@@
Despite these artifacts, Figure 9 quantitatively demonstrates that PSPC-Flex samples are still significantly more similar to network denoiser samples than the samples produced by the optimal or Gaussian denoisers. We measure similarity using the cosine similarity of SSCD descriptors (Pizzi et al., 2022), a embedding model used to detect image copies (details in Section C.1). Notably, across all network architectures, PSPC-Flex similarity scores approach 0.6, the threshold used to determine copying by Zhang et al. (2023).
Figure 9: SSCD cosine similarity of CIFAR-10 PF-ODE samples produced with varying denoisers from shared $\mathbf{z}$ initial conditions. Average PSPC-Flex similarity vs. neural-networks ($\mu=0.55$) is significantly higher than the similarity of optimal ($\mu=0.34)$ or Gaussian denoisers ($\mu=0.34$).
@@ 7 Related Work
Diffusion generalization Diffusion generalization has been considered in literature. Comparing a variety of networks, Zhang et al. (2023) finds diffusion models produce consistent samples despite differences in architecture and training. Both Niedoba et al. (2024) and Xu et al. (2023) find that diffusion models only deviate from empirical denoisers for intermediate $t$. Further, Niedoba et al. (2024) identify that errors in this region are primarily responsible for diffusion generalization. Similarly, Yi et al. (2023) attributes diffusion generalization to “slight differences” between the network and empirical denoisers. Kadkhodaie et al. (2024) suggest that generalization stems from geometrically adaptive harmonic bases. However, they do not consider how trained models deviate from empirical denoisers.
Several other methods have been proposed to reproduce diffusion generalization. Both Wang & Vastola (2024) and Li et al. (2024) find strong correlations between network denoisers and optimal denoisers under the simplistic model of a Gaussian $p(\mathbf{x})$. Closed-form diffusion models (Scarvelis et al., 2023) bias the empirical denoiser with a spectral bias (Rahaman et al., 2019) to produce a generalization mechanism. Concurrent to our work, Kamb & Ganguli (2024) use empirical patch denoisers to approximate diffusion model behaviour. However, their method uses square patches only, and uses a mixture of shared and localized patch collections.
Patch Denoising Patch based methods have long been used in the context of image denoising. For example, Field of Experts (Roth & Black, 2005) models images as Markov Random Fields to learn a prior over patches. Using expected patch log likelihood (Zoran & Weiss, 2011) and half quadratic splitting (Elad & Aharon, 2006; Zoran & Weiss, 2011; Friedman & Weiss, 2021), patch priors can be used to produce maximum a priori solutions to denoising problems. Additionally, Ding et al. (2023) and Wang et al. (2023) have recently proposed using patch-based approaches to improve diffusion model training.
@@ 8 Conclusions
Our work investigates the generalization mechanisms of image diffusion models. By comparing network denoisers and the optimal denoiser, we find strong evidence that a persistent local inductive bias in network denoisers results in denoiser approximation errors which are both qualitatively and quantitatively similar across all evaluated networks. We hypothesize that this local bias is the result of network denoisers employing local denoising operations to partially approximate the optimal denoiser. Approximating these local operations with patch posterior means, we find that these denoisers are excellent approximators of optimal denoiser patches over the majority of the forward diffusion process.
By spatially aggregating local patch denoisers, we find the resulting PSPC denoisers are remarkably similar to network denoisers when evaluated at the same $(\mathbf{z},t)$. Additionally, PSPC samples share structural similarity to those produced by diffusion models when sampling is identically initialized. We believe that the performance of PSPC is strong empirical evidence that a significant portion of the generalization behaviour of image diffusion models arises from a mechanism which is substantially similar to simple time-varying patch denoising and compositing operations.
Our work has numerous applications. Understanding the mechanisms of diffusion generalization helps to determine the cases when models fail to generalize, mitigating claims of “digital forgery” (Somepalli et al., 2023). Specifically, patch posterior probabilities (Equation 9) are a promising signal for identifying specific training images which influence the generation of a sample, with copyright and training-data licensing implications. In addition, patch-based diffusion generalization mechanisms may be exploited to improve training and sampling efficiency, as demonstrated by the preliminary exploration of Wang et al. (2023). Finally, further improvements to empirical denoisers such as PSPC-Flex may result in fully training-free models of similar quality to neural diffusion models, eliminating the fiscal and environmental costs of training.
@@ A.1 CIFAR-10
A summary of some of the differences between the networks compared in Figures 2 and 9 is given in Table 1
@@ C.1 SSCD Details
Appendix D Additional Denoiser Outputs
D.1 CIFAR-10
@@
Appendix E Additional PSPC-Flex Samples
We provide additional PSPC-Flex samples for CIFAR-10, FFHQ, and AFHQ datasets in Figures 21, 22 and 23. In each figure, each image in the top subplot has been generated from an identical latent seed from the corresponding image in the lower subplot.
Figure 21: Additional PSPC-Flex samples on CIFAR-10
Figure 22: Additional PSPC-Flex samples on FFHQ
Figure 23: Additional PSPC-Flex samples on AFHQ

### v3

@@ 5.1 PSPC-Square
A natural choice of patch set is $\mathcal{C}s$, the set of overlapping square patches of spatial size $s\times s$ defined in Section 4. However, Figure 4 shows that the optimal patch size varies substantially across the reverse diffusion process. To adapt our method to this observation, we define a patch size schedule $s(t):\mathbb{R}^{+}\rightarrow\{1,\ldots,h\}$. The square Patch Set Posterior Composite (PSPC-Square) is defined by Equation 9 with $\mathcal{C}=\mathcal{C}_{s(t)}$. In practice, we set s(t) according to Figure 4, choosing patch size at each $t$ with the lowest error versus the network denoiser. Patch size schedules can be found in Appendix B.
@@ 5.2 PSPC-Flex
While square patches mimic the square receptive fields of convolutional U-Nets, Figure 3 suggests that denoisers gradients are not perfectly square. To enable more flexible, non-square patches of varying shapes and sizes, we introduce PSPC-Flex, which utilizes adaptive patch sets based on average gradient sensitivity maps.
@@
Figure 7: Comparison of various denoisers against DDPM++ over forward and reverse processes. For every dataset, PSPC denoisers consistently provide better estimates of DDPM++ outputs than the optimal denoiser. In most cases, PSPC-Flex yields the best estimate of network denoisers, outperforming the Gaussian and CFDM denoisers. For comparisons against other network baselines, see Appendix D.
@@ 6.1 Network Denoiser Comparison
To evaluate the similarity of PSPC to network denoisers, we compare various empirical denoisers against the EDM DDPM++ denoiser outputs for each dataset. In addition to our PSPC-Square an PSPC-Flex methods, we evaluate the empirical denoiser of Equation 7, the Gaussian denoiser decribed by Wang & Vastola (2024); Li et al. (2024), and the Closed-Form Diffusion Model (CFDM) denoiser of Scarvelis et al. (2023).
Figure 7 plots the MSE of each of the denoisers against the neural network denoiser. We find that for both forward and reverse processes, both PSPC variants have substantially lower MSE against the network denoiser outputs as compared to the empirical and CFDM denoisers. Compared to the Gaussian denoiser, our methods generally have better MSE for $t\in[0.3,3]$. Comparing our methods, the flexibility of the PSPC-Flex patch set produces better estimates than those of PSPC-Square. Over the reverse process evaluation sets on every dataset, PSPC-Flex generally estimates the network denoiser better than PSPC-Square and all other methods.
Qualitatively, we compare the output of PSPC-Flex to several CIFAR-10 network denoisers and the empirical denoiser in Figure 1. For every $t$, PSPC-Flex outputs are visually similar to the outputs of the various network denoisers, suggesting that local denoising operations comprise a significant portion of the generalization mechanism of diffusion models. Additional denoiser outputs can be found in Appendix E.
Figure 8: PF-ODE sampling trajectories starting from the same initial $\mathbf{z}$, comparing the DDPM++ score estimator (Left) with our PSPC-Flex estimator (Right). At each $t$, we display the noisy image $\mathbf{z}$ along with the output of the DDPM++ denoiser and PSPC-Square on that $\mathbf{z}$. In all cases, minor differences between DDPM++ and PSPC-Square outputs only occur over intermediate $t$. Sampling using PSPC-Flex yields samples which are similar in content to the network samples. However, compounding errors lead to suboptimal final samples.
@@ 6.2 PSPC Sampling
Inspired by the remarkable similarity between network denoisers and PSPC-Flex evidenced by Figures 1 and 7, we investigate the efficacy of PSPC-Flex as fully training-free, completely non-neural diffusion model.
@@
Despite these artifacts, Figure 9 quantitatively demonstrates that PSPC-Flex samples are still significantly more similar to network denoiser samples than the samples produced by the optimal, CFDM, or Gaussian denoisers. We measure similarity using the cosine similarity of SSCD descriptors (Pizzi et al., 2022), a embedding model used to detect image copies (details in Section C.1). Notably, across all network architectures, PSPC-Flex similarity scores approach 0.6, the threshold used to determine copying by Zhang et al. (2023).
Figure 9: SSCD cosine similarity of CIFAR-10 PF-ODE samples produced with varying denoisers from shared $\mathbf{z}$ initial conditions. Average PSPC-Flex similarity vs. neural-networks ($\mu=0.55$) is significantly higher than the similarity of optimal ($\mu=0.34)$, CFDM ($\mu=0.23)$, or Gaussian ($\mu=0.34$) denoisers.
@@ 7 Related Work
Diffusion generalization There is a broad literature considering diffusion generalization. Comparing a variety of networks, Zhang et al. (2023) finds diffusion models produce consistent samples despite differences in architecture and training. Both Niedoba et al. (2024) and Xu et al. (2023) find that diffusion models only deviate from empirical denoisers for intermediate $t$. Further, Niedoba et al. (2024) identify that errors in this region are primarily responsible for diffusion generalization. Similarly, Yi et al. (2023) attributes diffusion generalization to “slight differences” between the network and empirical denoisers. Kadkhodaie et al. (2024) suggest that generalization stems from geometrically adaptive harmonic bases. However, they do not consider how trained models deviate from empirical denoisers.
Several other methods have been proposed to reproduce diffusion generalization. Both Wang & Vastola (2024) and Li et al. (2024) find strong correlations between network denoisers and optimal denoisers under the simplistic model of a Gaussian $p(\mathbf{x})$. Closed-form diffusion models (Scarvelis et al., 2023) bias the empirical denoiser with a spectral bias (Rahaman et al., 2019) to produce a generalization mechanism.
Concurrent to our work, Kamb & Ganguli (2024) propose a empirical, patch-based denoiser which leverages empirical patch posterior means and an increasing schedule of square patch sizes to approximate network denoiser outputs. While their method shares similarities with PSPC-Square, our work differs in several key ways. First, PSPC estimates patch-posterior means exclusively over localized patch sets. We found this to be particularly beneficial on datasets such as FFHQ, where image features are correlated to specific image positions. Second, while Kamb & Ganguli (2024) focus their analysis on convolutional diffusion denoisers, our work extends to both U-Net and DiT architectures. This broader evaluation is essential to our conclusion that local inductive biases form an important component of diffusion model generalization, regardless of architecture.
Patch-based Methods Patch-based methods have long been used in image processing. Classical image denoising methods such as Field of Experts (Roth & Black, 2005) model images as Markov Random Fields to learn patch-based priors. Methods such as expected patch log likelihood (Zoran & Weiss, 2011) and half quadratic splitting (Elad & Aharon, 2006; Zoran & Weiss, 2011; Friedman & Weiss, 2021) leverage such patch-based priors to enable maximum a priori image reconstructions. Importantly, these methods typically focus on recovering clean images rather than estimating posterior means, and typically operate on lower levels of additive Gaussian noise than is typical in the diffusion setting.
Patch-based methods have also been used in image generation. classical single image generation methods synthesize images by ensuring similarity between patches of the input and output images (De Bonet, 1997; Efros & Leung, 1999; Barnes et al., 2009; Simakov et al., 2008) while more recent, deep learning approaches leverage patches to train GANs(Shaham et al., 2019) and diffusion models (Nikankin et al., 2023; Wang et al., 2025). In the area of unconditional image generation, Ding et al. (2023) and Wang et al. (2023) utilize patch-based approaches to improve diffusion model training and performance.
@@ 8 Conclusions
Our work investigates the generalization mechanisms of image diffusion models. By comparing network denoisers and the optimal denoiser, we find strong evidence that a persistent local inductive bias in network denoisers results in denoiser approximation errors which are both qualitatively and quantitatively similar across all evaluated networks. We hypothesize that this local bias is the result of network denoisers employing local denoising operations to partially approximate the optimal denoiser. Approximating such local operations with patch posterior means, we find that these means are excellent approximators of optimal denoiser patches over the majority of the forward diffusion process.
By spatially aggregating local patch denoisers, we find the resulting PSPC denoisers are remarkably similar to network denoisers when evaluated at identical $(\mathbf{z},t)$ inputs. Additionally, PSPC samples share structural similarity to those produced by diffusion models when sampling is identically initialized. We believe that the performance of PSPC is strong empirical evidence that a significant portion of the generalization behaviour of image diffusion models arises from a mechanism which is substantially similar to simple time-varying patch denoising and compositing operations.
Our work has numerous applications. Understanding the mechanisms of diffusion generalization helps to determine the cases when models fail to generalize, mitigating claims of “digital forgery” (Somepalli et al., 2023). Specifically, patch posterior probabilities (Equation 9) are a promising signal for identifying specific training images which influence the generation of a sample, with copyright and training-data licensing implications. In addition, patch-based diffusion generalization mechanisms may be exploited to improve training and sampling efficiency, as demonstrated by the preliminary exploration of Wang et al. (2023). Finally, further improvements to empirical denoisers such as PSPC-Flex may result in fully training-free models of similar quality to neural diffusion models, eliminating the fiscal and environmental costs of diffusion training.
Acknowledgements
We acknowledge the support of the Natural Sciences and Engineering Research Council of Canada (NSERC), the Canada CIFAR AI Chairs Program, Inverted AI, MITACS, and Google. This research was enabled in part by technical support and computational resources provided by the Digital Research Alliance of Canada Compute Canada (alliancecan.ca), the Advanced Research Computing at the University of British Columbia (arc.ubc.ca), and Amazon.
@@ A.1 CIFAR-10
Our work analyzes diffusion generalization across a diverse set of network architectures, A summary of some of the differences between the networks compared in Figures 2 and 9 is given in Table 1
@@
In general, the diffusion models analyzed in this work produce high quality samples, as measured by Frechet Inception Distance (FID) (Heusel et al., 2017). The 50k FID scores for each model are given below in Table 3
Table 3: FID Scores for CIFAR-10 diffusion models
Model FID
DDPM++ (Karras et al., 2022) 1.97
NCSN++ (Karras et al., 2022) 1.98
DiT++ (Peebles & Xie, 2023) 9.08
U-ViT (Bao et al., 2022) 3.08
@@ C.1 SSCD Details
Appendix D Additional Network Baselines
Figure 2 compares four network denoisers on CIFAR-10 and finds that their outputs are qualitatively similar, and that they have similar mean squared error when compared against the optimal denoiser. Due to the similarity between networks, for clarity of presenation we utilized DDPM++ in Figures 3, 4 and 7 and did not present additional results using NCSN++, DiT or U-ViT network denoisers. For completeness, we have included additional versions of these figures below
D.1 Gradient Concentration
Figure 13 plots the average side length of a square patch required to capture a fixed percentage of the gradient concentration heatmap for DDPM++, NCSN++ and DiT network denoisers. Although there are some differences in the gradient concentrations, the general trend is consistent that network denoisers exhibit a more locally concentrated gradient for small $t$, and less concentration as $t$ increases.
Figure 13: Comparison of the Gradient concentration of DDPM++, NCSN++ and DiT networks across varying diffusion time. All three network denoisers display the same anti-correlation between gradient concentration and diffusion time.
D.2 Patch Errors
Figure 4 compares patch posterior means to patches of DDPM++ denoiser outputs for $\mathbf{z}$ drawn from that network’s reverse process. In Figure 14, we produce similar plots for DDPM++, NCSN++, DiT and U-ViT network denoisers. Across each denoiser, patch posterior means have similar errors when compared to patches of the network output. At both small and large $t$, appropriately sized patch posterior means have low MSE against network denoiser patches, with poorer MSE over intermediate $t$. Additionally, for each network we observe a similar trend that as $t$ increases, network denoiser patches are best estimated by patch posterior means of increasing spatial size.
Notably, for U-ViT, we find that patch posterior means do not estimate high $t$ denoiser outputs well. This is a symptom of the amplification problem for $\epsilon$-predictor networks, described by (Karras et al., 2022).
Figure 14: Comparison of patch posterior means of varying size against patches of network denoiser outputs for shared $\mathbf{z}$ drawn from each network’s reverse process on CIFAR-10. Across each network, we observe that patch posterior means are better estimators of network output patches than the optimal denoiser.
D.3 Denoiser Errors
Figure 7 visualizes the mean squared error between a variety of proposed denoisers and a DDPM++ network denoiser across the forward and reverse processes of three different datasets. However, other choices of network denoiser baseline are possible. Figure 15 plots denoiser MSE against four different network baselines. As checkpoints are only available on CIFAR-10 for DiT and U-ViT, no comparison is performed on FFHQ and AFHQ for these baselines. In addition to the denoisers shown in Figure 7, Figure 15 also plots MSE between pairs of network denoisers.
Examining Figure 15, PSPC-Flex consistently has lower MSE versus each network baseline than any other non-network denoiser. However, our method is consistently outperformed by the outputs of other network denoisers. We note that the gap between our approach and the performance of DiT when compared against DDPM++ and NCSN++ baselines is remarkably close. We believe closing this gap is a promising area of future research.
Figure 15: Comparison of network and empirical denoisers against varying network denoiser baselines. For DiT and UViT, no comparison is performed for FFHQ and AFHQ. Against each baseline, PSPC-Flex consistently outperforms Gaussian, CFDM, and optimal denoisers. However, PSPC performance consistently lags behind neural network performance.
Appendix E Additional Denoiser Outputs
E.1 CIFAR-10
@@
Appendix F Additional PSPC-Flex Samples
We provide additional PSPC-Flex samples for CIFAR-10, FFHQ, and AFHQ datasets in Figure 24. For each row, the images in the left subplot has been generated from an identical latent seed from the corresponding image in the right subplot. In addition, Figure 25 compares samples from a larger set of denoisers, including PSPC-Flex, PSPC-Square, NCSN++, Gaussian, optimal, and CFDM.
Figure 24: Additional PSPC-Flex PF-ODE samples on CIFAR-10, FFHQ and AFHQ. Samples in the left column are generated by PSPC-Flex, while those in the right column are generated by a DDPM++ network denoiser, starting from the same initial $\mathbf{z}$.
Figure 25: Comparison of PF-ODE CIFAR-10 samples generated using varying denoisers, starting from a shared initial $\mathbf{z}$.